OpenStudy (anonymous):

Factor 5a^4 – 5a completely. A. 5a(a – 1)(a^2 + a + 1) B. 5a(a^3) C. 5a(a + 1)(a^2 – a + 1) D. 5a(a – 1)(a + 1)

OpenStudy (souvik):

take 5a out..

OpenStudy (anonymous):

5a(a^3) ?

OpenStudy (souvik):

5a(a^3-1)

OpenStudy (anonymous):

But that is not in the answer choices!

OpenStudy (souvik):

now use..\[a ^{3}-b ^{3}=(a-b)(a ^{2}+ab+b ^{2)}\]

OpenStudy (anonymous):

sorry I don't understand:(

OpenStudy (souvik):

(a^3-1)=(a-1)(a^2+ab+b^2)

OpenStudy (souvik):

so its A